Find three consecutive odd integers such that their sum deceased by the second equals 50.
Equation-_________
Answer-__________

Found 2 solutions by addingup, rothauserc:
Answer by addingup(3677)

x+(x+2)+(x+2+2)-(x+2)= 50 Subtract the x+2s
x+(x+2+2)= 50
x+x+2+2= 50 Add on left
2x+4= 50 Subtract 4 both sides
2x = 46 Divide both sides by 2
x= 23
Proof:
23+25+27-25= 50
23+27= 50
50= 50

Answer by rothauserc(4718)

let x, x+2, x+4 be three consecutive odd integers, then
*******************************************************
Equation
x +x+2 +x+4 -(x+2) = 50
*******************************************************
Answer
3x +6 -x -2 = 50
2x +4 = 50
2x = 46
x = 23
therefore
23, 25, 27 is our answer
******************************************************
check answer
23 +25 +27 -25 = 50
23 + 27 = 50
50 = 50
our answer checks